### v1

# Real-time Event Recognition of Long-distance Distributed Vibration Sensing with Knowledge Distillation and Hardware Acceleration

Abstract
Distributed optical fiber vibration sensing (DVS) technology based on phase-sensitive optical time-domain reflectometry is widely used for safety monitoring and intrusion event surveillance in wide-ranging fields. Existing methods rely on deep learning models for event recognition but struggle with real-time processing of large data volumes in long-distance applications. To address these challenges, we use a four-layer convolutional neural network (CNN). The application of knowledge distillation with ResNet as the teacher model improves the generalization ability of the four-layer CNN, increasing the accuracy from 83.41% to 95.39% on data from untrained environments. The model is implemented on a field programmable gate array (FPGA) using a novel design that replaces multiplication with binary shift operations and quantizes model weights accordingly, allowing for high parallelism and low latency. An inference time of 0.083 ms is achieved for a spatial-temporal sample with a 12.5 m fiber length and 0.256 s time frame. This implies the system can process signals over a fiber length of approximately 38.55 km in real time, which is more than twice the capability of a GPU of Nvidia GTX 4090. The proposed method greatly improves the efficiency of vibration pattern recognition, thus promoting the application of DVS as smart sensing system in various areas. The data and code is available at https://github.com/HUST-IOF/Efficient-DVS.
Index Terms:
distributed vibration sensing, convolutional neural network, hardware acceleration, real-time system, fiber-optic IoT
I Introduction
Distributed optical fiber vibration sensing (DVS) technology is based on phase-sensitive optical time-domain reflectometry (phi-OTDR) system, and uses optical fibers as the sensing element. It is capable of providing dense and continuous wide-range measurement of vibration from surrounding environment and recognize various intrusion events based the sampled data. This technology has been applied as smart sensing system in various areas in recent years, including railway safety monitor[1, 2], perimeter security[3, 4], pipeline monitor[5, 6], and smart city[7]. As shown in Fig. 1, the phi-OTDR system captures the backscattering Rayleigh light across the fiber. The collected Rayleigh light trace, whose phase can be affected by external vibration, provides valuable information about the vibration events.
Figure 1: Schematic of DVS system. NLL: narrow linewidth laser, SOA: semiconductor optical amplifier, AWG: arbitrary waveform generator, EDFA: erbium-doped fiber amplifier, FUT: fiber under test, APD: avalanche photodetector, DAQ: data acquisition card.
Signal processing algorithms are able to extract information from vibration traces. However, the recognition accuracy of vibration events is often limited by the variability of signals caused by complex and changing environments, as well as the diverse types of intrusion events encountered in practical applications. To improve the recognition performance, conventional methods typically involve hand-crafted algorithms for data preparation and feature extraction, followed by the use of classic machine learning algorithms for recognition[8, 9, 10, 11, 12, 13]. However, this kind of approaches can be time-consuming in both development and deployment. Designing feature extraction algorithms requires researchers to possess knowledge of signal processing algorithms and the specific application context. As a result, the resulting algorithms can be complex, time-consuming, and may not capture the most important features accurately.
To address these challenges, deep learning models (DLMs) have been introduced to the field. Initially, DLMs were introduced for pattern recognition based on extracted features [9]. With the development of more sophisticated and computationally powerful DLMs, they are capable of automatically extracting important features from noisy data and simultaneously handling feature extraction and pattern recognition tasks. The introduction of complex DLMs brings about improved performance and reduces the design workload. Consequently, they have received more attention in recent years. Table I shows several typical methods of this type of application, where $GAF$ represents gramian angular field, $S-T$ denotes space-time, and $F-T$ is frequency-time.
TABLE I: Previous Work on DLM-based DVS Algorithms and Our Work
Model Input Form Parameter Number Accuracy (%) Pub. Year
VGG16 GAF Figure 138M 97.67 2020[3]
YoloV3 S-T Image 45.5M - 2021[14]
Yolo-A30 S-T Image 65M 99.2 2022[2]
Yolo S-T Image 61M 96.1 2022[15]
MTL S-T Image 1.13M 99.46 2023[7]
Swin-T GAF Figure 28M 92.44 2023[6]
Resnet-152 F-T Image 60M 96.67 2023[16]
CNN-4 (This work) S-T Image 30K 95.39 -
However, the utilization of high-capacity DLMs has resulted in increased computation resource requirements. Moreover, when using the DVS system to monitor optical fibers, it is important to note that the volume of captured data directly relates to the length of the fiber. Due to the limited available computation resources in real-world applications, achieving real-time recognition is only possible over short-distance fibers [17]. To improve the computational efficiency of the DVS system, there are typically three solutions. The first is data selection or down-sampling, which involves using prior knowledge and data analysis to identify and retain valuable signal sections or sample points[2, 7]. This approach aims to reduce the input data volume through hand-crafted algorithms, though it may result in false negatives. The second solution is model compression, which seeks to lower computational complexity and resource demands. However, existing solutions use models with a parameter number no lower than millions to ensure performance and generalizability[7]. So, there is still potential for further improvement to create smaller and more efficient models. The third solution is the use of high-performance hardware, such as graphics processing units (GPUs) and cloud computing. While this approach can significantly increase computational power, it is often expensive and demands high power consumption, as well as a reliable communication system to handle the large volume of sample data.
This paper proposes to improve the signal processing capability of the DVS system through knowledge distillation and hardware acceleration. We propose to use a shallow convolutional neural network (CNN), instead of DLMs, to reduce model complexity and computation resource requirements. Moreover, we identify the generalizability limitation of low-capacity models and propose to use the knowledge distillation technique as a solution. To further improve the throughput of the data processing system, we design an efficient hardware acceleration scheme based on field programmable gate array (FPGA). Additionally, we have integrated an Analog-to-Digital Converter (ADC) with the FPGA to minimize latency during data transfer, as depicted in Fig. 1. The proposed scheme utilizes FPGA’s flexibility to achieve full optimization of the model. It involves replacing multiplication operations in the model inference with shift operations, as the later ones can be implemented without digital signal processing (DSP) resources. This design is only limited by the logic resources, instead of DSP resources which is more limited[18], on the chip, allows for higher parallelism and thus faster calculation. This architecture is referred as the shift-add structure in the rest of the paper. Additionally, we propose to quantize the model weights into integers that indicate the number of binary shift operations. The quantization technique is referred as shift parameter quantization in the rest of the paper. Through this approach, we can achieve real-time processing of a large amount of captured data from the DVS system on a mainstream industrial FPGA.
In a word, the contribution of this work can be listed as follows:
  • We propose a novel scheme for designing and implementing real-time DVS algorithms and evaluate it using real-world DVS datasets. The results demonstrate comparable performance and generalizability to deep learning models, while achieving real-time recognition over long-distance fibers.
  • Shallow CNN models, coupled with knowledge distillation, are used to develop a lightweight model that offers high performance and generalizability. This approach effectively addresses the throughput limitation brought by the high computational complexity of existing algorithms.
  • A novel hardware acceleration scheme targeted for shallow CNN is developed, leveraging shift-add structures and shift parameter quantization to achieve higher efficiency.
II Method
II-A Knowledge Distillation based on logits
Figure 2: Schematic of KD process.
It is generally believed that large models, that is, models with more complex and efficient structural design and huge number of parameters, tend to be able to extract higher-level features, so that it can achieve better results in various tasks. In other words, large models often have better generalizability, and can better handle signals with various noise. On the other hand, the model inference process can be considered as finishing the specified task based on the ’knowledge’ contained in model structure and its parameters. The idea of knowledge distillation is to extract the knowledge from large models (teacher models), and transfer the knowledge to small models (student model), so that the performance and generalizability of the later one can be improved. In the knowledge distillation method based on logits[19], the output prediction distribution of the model is considered to contain valuable knowledge. This distribution represents the probabilities of classifying the input sample into each class. The distribution can be obtained through introducing a temperature parameter $T$ to the softmax function at output layer. The function can be expressed as:
$\text{Softmax}(\mathbf{x}/T)_{i}=\frac{e^{x_{i}/T}}{\sum_{j}e^{x_{j}/T}}$ (1)
where $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]$ is the input vector, which is the output of the network in this case. The knowledge is then transferred by guiding the student model to imitate the behavior of the teacher model. The process is to minimizing a loss function that measures the difference between the outputs of the teacher network and the student network, and also the difference between the outputs of the student network and the correct results, as shown in Fig. 2. The process can be seen as minimizing the difference between the outputs of the student network and the correct results. It can be achieved through minimizing a loss function that measures the difference between the outputs of the teacher network and the student network. The loss function usually consists of a standard classification loss, like cross-entropy, and a distillation loss that measures the difference between the softened probabilities generated by the teacher and student networks. Formally, the loss function $\mathcal{L}$ can be expressed as:
$\begin{split}\mathcal{L}&=\alpha\cdot\text{CE}(\mathbf{y}_{s},\mathbf{y}_{\text{true}})\\ &\quad+(1-\alpha)\cdot\text{KL}(\text{Softmax}(\mathbf{y}_{t}/T),\text{Softmax}(\mathbf{y}_{s}/T))\end{split}$ (2)
where CE denotes the cross-entropy loss, KL represents the Kullback-Leibler divergence, $\mathbf{y}_{\text{true}}$ is the true label, $\alpha$ represents wegiht between two kind of losses, $\mathbf{y}_{t}$ and $\mathbf{y}_{s}$ denote the outputs of the teacher and student networks respectively. Through this optimization process, knowledge distillation enables the creation of compact yet powerful neural networks capable of retaining the performance of their larger counterparts.
II-B FPGA-Based Hardware Acceleration
II-B1 Parallel Architecture
The main problem of implementing CNN on FPGA is the storage of intermediate data generated during the calculation process and the parameters of the model[20]. There are two typical solutions to the problem. One option is to use external memory to expand the storage capacity[21, 22, 23, 18]. This type of design usually combines on-chip storage and external storage to form a cache system. The on-chip computing unit and memory access pattern is designed in accordance with the bandwidth limitation of the cache system. The model is divided into multiple computing tasks suitable for the computing unit. As the tasks are completed in sequence, the calculation of the entire model is also completed. Another option is to compress the model and store all intermediate data and parameters on-chip[20, 18]. Such design typically adopts a pipeline structure, implementing the model as multiple parallel modules to improve the throughput of the design. In this paper, in order to achieve faster inference efficiency, the later one is adopted. Each layer of the model is implemented as a separate module as shown in Fig. 3.
Figure 3: Schematic of pipelined structure.
II-B2 Activation Buffering
The activation buffer, as shown in Fig. 3, stores intermediate data from the preceding layer. A typical convolution and max pooling layer can be expressed as follows:
$\begin{split}O_{conv}[m,x,y]&=\sum_{n=0}^{N-1}\sum_{p=0}^{P-1}\sum_{q=0}^{Q-1}I[n,x+p,y+q]\\ &\quad*K[m,n,p,q]+B[m]\\ \end{split}$ (3)
$O_{max}[n,x,y]=\max\{I[n,x:x+P,y:y+Q]\}$ (4)
Here, the symbols $O$, $I$, $M$, $N$, $P$, $Q$, $K$, and $B$ represent the following components: $O$ denotes the output feature map, $I$ represents the input feature map, $M$ corresponds to the number of output channels, $N$ denotes the number of input channels, $P$ represents the kernel height, $Q$ represents the kernel width, $K$ symbolizes the convolution kernel, and $B$ signifies the bias term.
Obviously, both convolution layers and pooling layers only require the input feature map data within the selected window. Assume the input to be a spatial-temporal figure, where rows represent discrete times and columns represent positions. The data is streamed in row-major order. Then the activation buffer only needs to store $P$ rows from the input feature map during the computation process. Thus, the buffer mechanism is designed as shown in Fig. 4.
Figure 4: Line buffer structure.
When new data is inputted, the corresponding column shifts the existing data up by one unit, and the new data is filled into the bottom unit. This design allows the buffer to store just $P-1$ rows and $Q$ data from the last row to start the convolution operation. Thus, the design only requires $P$ rows of storage capacity. Moreover, reading new data and calculation can be performed in parallel to achieve lower latency. The convolution layer only needs to buffer $(P-1-S)*W+Q$ data for each channel instead of the complete input feature map to start operation, where $W$and $S$ denote the width of input feature map and the stride respectively. Thank you for the clarification. I apologize for the misunderstanding. In the case of a convolution layer, it only needs to buffer $(P-1-S)\times W+Q$ data for each channel, rather than the entire input feature map, in order to commence the operation. Here, $W$ represents the width of the input feature map, and $S$ represents the stride.
II-B3 Shift-Add Structure
The main operation of convolutional neural networks is multiplication and accumulation (MAC). Digital signal processing (DSP) block is required to implement MAC units on FPGA. The limited number of DSP blocks restricts the number of MAC units that can be implemented on-chip. As a result, the throughput of the design is restricted. To avoid this restriction, it is proposed to use the shift and accumulation units that do not require DSP resources for implementation to replace the MAC units, as shown in Fig. 5. Digital values can be represented as a sum of powers of 2, as illustrated in Equation 5 with exponents denoted as $s_{0},s_{1},\ldots,s_{k}$. Multiplication with power of two is essentially performing shift operations on multiplier. This allows a MAC operations to be losslessly replaced by a series of shift and accumulation operations, as shown in Equation 6.
Figure 5: Shift-add and MAC structure.
$W=2^{s_{0}}+2^{s_{1}}+\cdots+2^{s_{k}}$ (5)
$\begin{split}I\cdot W&=I\cdot(2^{s_{0}}+2^{s_{1}}+\cdots+2^{s_{k}})\\ &=I<<s_{0}+I<<s_{1}+\cdots+I<<s_{k}\end{split}$ (6)
II-B4 Shift Parameter Quantization
After converting all the MAC operations to shift and accumulation operations, the powers of two in the above equation 5 and 6 can be seen as separate parameters representing number of shift and accumulation operations, which are referred as shift parameters in this paper. To reduce the latency and demand for storage space of the design, a magnitude-based post-training pruning strategy is employed here to reduce the number of the shift parameters[24]. It is to assume that the smaller parameters have less impact on the final output, i.e., less significant. Based on this assumption, the smaller parameters can be discarded without severe influence on the final performance of the model. The complete quantization process is briefly described in detail with pseudo code shown in Algorithm 1.
Algorithm 1 Shift Parameter Quantization
1:Input
2:     $Params$matrix of parameters
3:     $N$Number of parameters remained
4:Output
5:     $Shifts$Quantized parameters
6:     $Signs$Sign of parameters
7:
8:$Signs\leftarrow\textsc{sign}(Params)$
9:$[rows,cols]\leftarrow\textsc{shape}(Params)$
10:for $i=1$ to $rows-1$ do
11:     for $j=1$ to $cols-1$ do
12:         $BitPositions\leftarrow\textsc{zeros}(N)$
13:         $count\leftarrow 0$
14:         $P_{binary}\leftarrow\textsc{binary}(AbsParams[i][j])$
15:         for $i\leftarrow 0$ to $\textsc{len}(P_{binary})-1$ do
16:              if $count<N$ then
17:                  if $P_{binary}[i]=1$ then
18:                       $BitPositions[count]\leftarrow i+1$
19:                       $count\leftarrow count+1$
20:                  else
21:                       $BitPositions[count]\leftarrow None$
22:                  end if
23:              end if
24:         end for
25:         $Shifts[i][j][:]\leftarrow BitPositions$
26:     end for
27:end for
Conventional quantization methods typically involve mapping weights to a limited integer range[25]. The output data of the quantized layer requires to be mapped back to the original value range before it can be used as input of the subsequent layers. The proposed method does not require such operations. Thus, the hardware design does not need any DSP block to implement the floating-point multipliers. As a result, the design can avoid potential throughput limitations related to DSP resources limitation.
II-B5 Encoding
After quantization, the next step is to encode the parameters for storage. In order to further compress the storage space required for parameters, offset binary encoding is used[26]. The offset binary encoding scheme introduces an offset to reposition the data within an optimal range before performing binary encoding. In this scheme, the data is moved towards zero. In the encoding process, we first extract and save the sign information of the parameters. Then, we perform quantization on the absolute value of the parameters. The bias is chosen based on the smallest non-zero number to ensure that it is moved to zero. The parameters that are outside the data range of the majority of parameters are rounded to the nearest number. The pseudo code of the process is shown in Algorithm 2.
Algorithm 2 Biased Binary Encoding
1:Input
2:     $Params$matrix of parameters
3:     $N$Number of bits used for encoding
4:Output
5:     $Encoded$Encoded parameters
6:     $Bias$Bias for the encoding
7:     $Signs$Sign of parameters
8:
9:$MinValue\leftarrow\textsc{min}(Params)$
10:$Bias\leftarrow-MinValue$
11:$MaxRange\leftarrow 2^{N}-1$
12:$Signs\leftarrow\textsc{sign}(Params)$
13:$AbsParams\leftarrow\textsc{abs}(Params)$
14:$[rows,cols]\leftarrow\textsc{shape}(Params)$
15:for $i=1$ to $rows-1$ do
16:     for $j=1$ to $cols-1$ do
17:         $P_{biased}\leftarrow AbsParams[i][j]+bias$
18:         if $P_{biased}>MaxRange$ then
19:              $P_{biased}\leftarrow MaxRange$
20:         end if
21:         $P_{encoded}\leftarrow\textsc{binary}(P_{biased})$
22:         $Encoded[i][j]\leftarrow P_{encoded}$
23:     end for
24:end for
III Experiment
The proposed scheme is evaluated using a real-world dataset captured by a DVS system, as illustrated in Fig. 1. The system capture a trace of Rayleigh back scattering light every 1 ms, with a spatial interval of 1.25 m. The samples stored as a 2-dimensional matrix with a size of 256 by 11, represent spatial-temporal figures, capturing a segment of the fiber with a length of 12.5 m and a time frame of 0.256 s. A 30 km long fiber was utilized, with a section of 50 m buried under 0.5 m of composite material composed of soil, sand, and stones in random ratios. The composition of the material has a significant impact on the vibration signal detected by the fiber, making it an effective means to explore the influence of environmental complexity. The dataset comprises two parts, collected from different locations at different time with varying material compositions, simulating the installation of the system in different environments for the same application. Specific details regarding the composition of the datasets are provided in Table II. Typical data are shown in Fig. 6.
Figure 6: Visualization of typical data samples. (a) Air Pick. (b) Excavator. (c) Hammer.
TABLE II: Class label and distribution of samples of the datasets
Class Dataset 1 Dataset 2
Hammer 3332 268
Air Pick 3558 330
Excavator 3359 277
III-A Performance evaluation
To evaluate and demonstrate the effectiveness of the proposed scheme, we employ a 5-fold cross-validation strategy on Dataset 1. This strategy involves dividing the dataset into five equally-sized subsets or folds. The division of folds remains consistent throughout the subsequent evaluations to ensure fair comparison of results. During each iteration of the cross-validation process, the model is trained on four folds while the remaining fold is held out for validation. This process is repeated five times, with each fold serving as the validation set once. The Dataset 2 is used as test set to verify the generalizability of the model. By averaging the performance of the five runs, we obtain the final validation and test result. Adam optimizer is used for training. Learning rate is set to 0.001, and it is halved when the training loss does not decrease within five epochs. The training setting is also kept consistent for the subsequent evaluations.
TABLE III: Parameters of the CNN Model
Layers Kernel Size Stride Padding Ouput Channels
Conv 1 3*3 1 1 8
MaxPool 1 2*2 2 0 8
Conv 2 3*3 1 1 16
MaxPool 2 2*2 2 0 16
Conv 3 3*3 1 1 32
AvgPool 1 2*2 2 0 32
Conv 4 3*3 1 1 64
flatten - - - 2048
FC 1 - - - 3
To evaluate and demonstrate the performance of the lightweight model, a 4-layer CNN with a structure outlined in Table III is utilized. The model consists of four layers and has a total of 30,771 parameters. In terms of computational cost, it requires 2,282,496 floating-point operations (FLOPs). Additionally, this 4-layer CNN model serves as the baseline for evaluating the improvements in both performance and generalizability achieved through knowledge distillation.
The baseline model achieves an average validation accuracy of 99.69%, while the average test accuracy is 83.41%. The feature visualization of the best model on the test set is shown in Fig. 7a. The drop in accuracy clearly demonstrates the limitation of generalizability of the baseline model.
TABLE IV: Evaluation Results of the Resnet Models
Model Val/Test/FLOPs Params
Resnet-18 99.58 95.47 98472960 2777283
Resnet-34 99.50 97.67 226235392 8164803
Resnet-101 98.34 85.87 857427968 27532227
Resnet-152 99.07 83.04 552325120 43175875
To ensure a strong baseline for knowledge distillation, we employed the Resnet model, which is widely used for this task in the relevant research[27, 28, 29], as the teacher model. We further evaluated Resnet models with varying structures and depths on the dataset, and the results are presented in Table IV. Considering our goal of improving both the performance and generalizability of the student model through knowledge distillation, we selected the Resnet-34 model as the teacher model. The knowledge distillation process introduces two additional hyper-parameters, namely $\alpha$ and $T$. To find the best combination, we iterate through different values of $\alpha$ from 0 to 1 with a step of 0.1, and $T$ from 1 to 10 with a step of 1. The optimal combination is found to be $\alpha=0.1$ and $T=5$.
Through knowledge distillation, the model achieves an average test accuracy of 95.39%, while maintaining an average validation accuracy of 99.61%. The feature visualization on the test set is shown in Fig. 7b, to better demonstrate the impact of knowledge distillation. It is clear that the model trained with knowledge distillation exhibits a more distinct boundary between samples belonging to different categories, demonstrating an improvement generalizability compared with the baseline model. The comparison of the baseline model, teacher model, and student model is presented in Fig. 8. According to the figure, it is evident that the baseline model, which has fewer capacity and depth compared to the teacher model, exhibits limitations in terms of generalizability. The baseline model fails to extract high-level features that are less influenced by the training dataset, resulting in poor and unstable performance on the test dataset. It also shows that the knowledge distillation technique enables a lightweight CNN model, with only 30771 parameters and requiring 2282496 FLOPs, to achieve generalizability comparable to a larger model with 8164803 parameters and 226235392 FLOPs.
Figure 7: Feature visualization of the models. (a) Baseline Model. (b) Improved Model.
Figure 8: Validation accuracy and test accuracy of the models.
III-B Implementation Evaluation
We conduct a performance evaluation of the proposed quantization method on the resulting KD-improved models. The entire model is quantized together, meaning that all weights are quantized to the same number of shift parameters. During the quantization process, we iterate through different values of the number of shift parameters, ranging from 1 to 10. This enables us to assess and analyze the performance of the model across different degrees of quantization, thus allowing us to identify the optimal balance between compression level and performance degradation. The results of the quantization process are presented in Fig. 9a, which shows the performance of the quantized models on the validation set. Based on the results, it can be determined that the optimal number of shift parameters to be used is 3. This allows the compressed model to achieve the same performance as the original model. To validate the generalizability of the model, the quantization process is further assessed on the test dataset, as demonstrated in Fig. 9b. The results clearly indicate that the compressed model performs comparably to the original model, without any noticeable degradation in generalizability. This choice of quantization configuration effectively preserves all the essential information of the model while achieving the desired performance on the test dataset.
Figure 9: Change in accuracy with number of shift parameters: (a) Accuracy change on the validation set; (b) Accuracy change on the test set.
The encoding process is then performed to optimize the on-chip storage. It is carried out on a layer-wise basis, which means that the parameters of the same layer are encoded together using the same bias. To determine the optimal number of bits used for encoding, we conduct iterations ranging from 1 to 8. The results obtained from the validation dataset, as depicted in Fig. 10a, indicate that employing a 3-bit encoding scheme achieves lossless compression. Furthermore, the encoding process is performed on the test dataset and visualized in Fig. 10b. It is evident that the encoding procedure brings a certain degree of performance degradation in terms of the generalizability of the model. Nevertheless, despite this influence, it is worth noting that the impact remains stable and minimal. Such stability in performance degradation implies that the encoding process does not significantly influence the overall performance of the model.
Figure 10: Change in accuracy with number of bits used for encoding: (a) Accuracy change on the validation set. (b) Accuracy change on the test set.
It is clear that the proposed approach offers an effective solution for achieving compression, with only a minimal impact on the generalizability of the model, as shown by the aforementioned results. We can then calculate the compression rate using the following information: the model initially consists of 30771 32-bit floating-point format parameters, and each weight is compressed into three 3-bit shift parameters. Consequently, the compression rate can be calculated: $3\times 3/32\times 100\%=28.125\%$. Additionally, the proposed compression scheme has been adapted to the shift-add structure, further enhancing the overall system performance.
TABLE V: Performance of the Models on Different Hardware
Model Inference time Inference time Inference time
on CPU/ms on GPU/ms on FPGA/ms
ResNet-34 1.846 0.375 -
Quantized CNN 0.254 0.178 0.083
The optimized model is then implemented with proposed FPGA-based hardware acceleration scheme. To evaluate the proposed scheme, we use the state-of-the-art CPU and GPU as the baseline performance. Here, the utilized CPU is Intel i9-14900, GPU is Nvidia GTX 4090. The average inference time of them is shown in Table V. The FPGA chip utilized here is Xilinx ZCU15eg. For quick development and fast implementation, the design is built with Xilinx high level HLS. The final design achieved a latency of 25112 cycles running at a frequency of 303 MHz. The inference time of the system can be calculated as $25112\times 1/(303\times 106)\approx 0.083ms$. The sample time for each data frame is 256 ms, allowing the system to process 3,084 frames within this interval. This processing capacity is equivalent to handling the data generated when the system monitors a fiber with a length of $12.5\times 3084=38550m$. Therefore, the system can achieve real-time processing over a 38-km fiber, which is longer than the 30-km fiber utilized for sampling the dataset. Based on bibliographical research, the achieved recognition time of 0.083 ms is the fastest reported so far in the field [2, 7].
TABLE VI: Hardware Resource Utilized
Resource Utilization Available
LUT 280828 341280
FF 257011 682560
DSP 0 3528
BRAM 11 744
URAM 0 112
The resource requirement of the design is shown in Table VI. Our design makes efficient use of on-chip resources like look-up tables (LUT) and flip-flops (FF), allowing for high parallelism without being limited by the amount of DSP resources available. This optimization enables our system to achieve a higher performance compared with the CPU and GPU solutions.
IV Conclusion
This paper addresses the throughput limitation issue faced by the DVS systems based on DLMs, which prevents the implementation of real-time recognition over long-distance fibers. To overcome this challenge, we propose to use lightweight models as a replacement for DLMs to reduce the computation resource requirements of the system. We identify the problem of limited generalizability of the lightweight models and adopts knowledge distillation techniques to enhance the generalizability of the utilized lightweight model. Furthermore, we propose an FPGA-based hardware acceleration scheme for fast inference. It results in an impressive 2.14x shorter inference time compared to the state-of-the-art GPU implementation, and 4.52x compared to the CPU implementation, which is also the fastest recognition time reported so far in this field. Through our proposed method, we achieve real-time recognition over 38-km-long fibers. The implemented system allows for real-time recognition over long-distance fibers in practical applications, contributing to the advancement of DVS technology and its use as a smart sensing system in various areas.

### v2

@@
 Abstract
-Distributed optical fiber vibration sensing (DVS) technology based on phase-sensitive optical time-domain reflectometry is widely used for safety monitoring and intrusion event surveillance in wide-ranging fields.
-Existing methods rely on deep learning models for event recognition but struggle with real-time processing of large data volumes in long-distance applications.
-To address these challenges, we use a four-layer convolutional neural network (CNN).
-The application of knowledge distillation with ResNet as the teacher model improves the generalization ability of the four-layer CNN, increasing the accuracy from 83.41% to 95.39% on data from untrained environments.
-The model is implemented on a field programmable gate array (FPGA) using a novel design that replaces multiplication with binary shift operations and quantizes model weights accordingly, allowing for high parallelism and low latency.
-An inference time of 0.083 ms is achieved for a spatial-temporal sample with a 12.5 m fiber length and 0.256 s time frame.
-This implies the system can process signals over a fiber length of approximately 38.55 km in real time, which is more than twice the capability of a GPU of Nvidia GTX 4090.
-The proposed method greatly improves the efficiency of vibration pattern recognition, thus promoting the application of DVS as smart sensing system in various areas.
-The data and code is available at https://github.com/HUST-IOF/Efficient-DVS.
+Fiber-optic sensing, especially distributed optical fiber vibration (DVS) sensing, is gaining importance in internet of things (IoT) applications, such as industrial safety monitoring and intrusion detection.
+Despite their wide application, existing post-processing methods that rely on deep learning models for event recognition in DVS systems face challenges with real-time processing of large sample data volumes, particularly in long-distance applications.
+To address this issue, we propose to use a four-layer convolutional neural network (CNN) with ResNet as the teacher model for knowledge distillation.
+This results in a significant improvement in accuracy, from 83.41% to 95.39%, on data from previously untrained environments.
+Additionally, we propose a novel hardware design based on field-programmable gate arrays (FPGA) to further accelerate model inference.
+This design replaces multiplication with binary shift operations and quantizes model weights, enabling high parallelism and low latency.
+Our implementation achieves an inference time of 0.083 ms for a spatial-temporal sample covering a 12.5 m fiber length and 0.256 s time frame.
+This performance enables real-time signal processing over approximately 38.55 km of fiber, about 2.14×2.14$\times$2.14 × the capability of an Nvidia GTX 4090 GPU.
+The proposed method greatly enhances the efficiency of vibration pattern recognition, promoting the use of DVS as a smart IoT system.
+The data and code are available at https://github.com/HUST-IOF/Efficient-DVS.
 Index Terms:
@@
 I Introduction
-Distributed optical fiber vibration sensing (DVS) technology is based on phase-sensitive optical time-domain reflectometry (phi-OTDR) system, and uses optical fibers as the sensing element.
-It is capable of providing dense and continuous wide-range measurement of vibration from surrounding environment and recognize various intrusion events based the sampled data.
-This technology has been applied as smart sensing system in various areas in recent years, including railway safety monitor[1, 2], perimeter security[3, 4], pipeline monitor[5, 6], and smart city[7].
+Distributed optical fiber vibration sensing (DVS) technology, based on phase-sensitive optical time-domain reflectometry (phi-OTDR), uses optical fibers as sensing elements to provide dense, continuous measurements of vibration and recognize various intrusion events.
+This technology has been applied as smart internet of things (IoT) systems in various areas in recent years, including industrial safety monitoring and intrusion detection.
+Recent applications of DVS technology include railway safety monitoring [1, 2], perimeter security [3, 4], pipeline monitoring [5, 6], and smart cities [7].
 As shown in Fig.
-1, the phi-OTDR system captures the backscattering Rayleigh light across the fiber.
-The collected Rayleigh light trace, whose phase can be affected by external vibration, provides valuable information about the vibration events.
+1, the phi-OTDR system captures Rayleigh backscattering light across the fiber, with phase variations induced by external vibrations providing valuable event information.
 Figure 1: Schematic of DVS system.
 NLL: narrow linewidth laser, SOA: semiconductor optical amplifier, AWG: arbitrary waveform generator, EDFA: erbium-doped fiber amplifier, FUT: fiber under test, APD: avalanche photodetector, DAQ: data acquisition card.
-Signal processing algorithms are able to extract information from vibration traces.
-However, the recognition accuracy of vibration events is often limited by the variability of signals caused by complex and changing environments, as well as the diverse types of intrusion events encountered in practical applications.
-To improve the recognition performance, conventional methods typically involve hand-crafted algorithms for data preparation and feature extraction, followed by the use of classic machine learning algorithms for recognition[8, 9, 10, 11, 12, 13].
-However, this kind of approaches can be time-consuming in both development and deployment.
-Designing feature extraction algorithms requires researchers to possess knowledge of signal processing algorithms and the specific application context.
-As a result, the resulting algorithms can be complex, time-consuming, and may not capture the most important features accurately.
-To address these challenges, deep learning models (DLMs) have been introduced to the field.
+Despite its capabilities, DVS technology faces several challenges in IoT applications.
+The primary issue is the variability of signals due to complex and changing environments and diverse intrusion types, which limits event recognition accuracy.
+Traditional methods often rely on hand-crafted algorithms for data preparation and feature extraction, followed by classic machine learning for recognition[8, 9, 10, 11, 12, 13].
+These approaches can be time-consuming and complex, requiring extensive expertise in signal processing.
+To solve these problems, deep learning models (DLMs) have been introduced to the field.
 Initially, DLMs were introduced for pattern recognition based on extracted features [9].
@@
 CNN-4 (This work) S-T Image 30K 95.39 -
-However, the utilization of high-capacity DLMs has resulted in increased computation resource requirements.
-Moreover, when using the DVS system to monitor optical fibers, it is important to note that the volume of captured data directly relates to the length of the fiber.
-Due to the limited available computation resources in real-world applications, achieving real-time recognition is only possible over short-distance fibers [17].
+Another challenge in applying DVS technology in IoT applications is the large volume of captured data, which scales with the length of the fiber.
+Consequently, monitoring long-distance fibers requires processing vast amounts of data.
+Moreover, applications such as pipeline monitoring and earthquake detection demand real-time or near-real-time responses from smart IoT systems based on DVS technology to minimize potential property damage.
+In addition, the use of high-capacity DLMs significantly increases the demand for computational resources.
+In real-world scenarios, where resources are often limited, achieving real-time event recognition is typically feasible only for short-distance fiber monitoring [17].
 To improve the computational efficiency of the DVS system, there are typically three solutions.
 The first is data selection or down-sampling, which involves using prior knowledge and data analysis to identify and retain valuable signal sections or sample points[2, 7].
-This approach aims to reduce the input data volume through hand-crafted algorithms, though it may result in false negatives.
+While this reduces the input data volume, it can lead to false negatives due to its reliance on hand-crafted algorithms.
 The second solution is model compression, which seeks to lower computational complexity and resource demands.
-However, existing solutions use models with a parameter number no lower than millions to ensure performance and generalizability[7].
-So, there is still potential for further improvement to create smaller and more efficient models.
+However, existing approaches often require models with at least millions of parameters to maintain performance and generalizability[7], leaving room for further optimization to develop smaller and more efficient models.
 The third solution is the use of high-performance hardware, such as graphics processing units (GPUs) and cloud computing.
 While this approach can significantly increase computational power, it is often expensive and demands high power consumption, as well as a reliable communication system to handle the large volume of sample data.
-This paper proposes to improve the signal processing capability of the DVS system through knowledge distillation and hardware acceleration.
-We propose to use a shallow convolutional neural network (CNN), instead of DLMs, to reduce model complexity and computation resource requirements.
-Moreover, we identify the generalizability limitation of low-capacity models and propose to use the knowledge distillation technique as a solution.
-To further improve the throughput of the data processing system, we design an efficient hardware acceleration scheme based on field programmable gate array (FPGA).
-Additionally, we have integrated an Analog-to-Digital Converter (ADC) with the FPGA to minimize latency during data transfer, as depicted in Fig.
+This paper proposes enhancing the signal processing capabilities of DVS systems through the use of shallow convolutional neural network (CNN) coupled with knowledge distillation and hardware acceleration.
+We introduce a shallow CNN as a more efficient alternative to DLMs, reducing model complexity and computational resource requirements.
+we identify the generalizability limitation of low-capacity models and propose to use the knowledge distillation technique as a solution.
+To further improve the data processing throughput, we design an efficient hardware acceleration scheme based on a field programmable gate array (FPGA).
+An analog-to-digital Converter (ADC) is integrated with the FPGA to minimize latency during data transfer, as depicted in Fig.
 1.
-The proposed scheme utilizes FPGA’s flexibility to achieve full optimization of the model.
-It involves replacing multiplication operations in the model inference with shift operations, as the later ones can be implemented without digital signal processing (DSP) resources.
-This design is only limited by the logic resources, instead of DSP resources which is more limited[18], on the chip, allows for higher parallelism and thus faster calculation.
-This architecture is referred as the shift-add structure in the rest of the paper.
-Additionally, we propose to quantize the model weights into integers that indicate the number of binary shift operations.
-The quantization technique is referred as shift parameter quantization in the rest of the paper.
-Through this approach, we can achieve real-time processing of a large amount of captured data from the DVS system on a mainstream industrial FPGA.
+The proposed scheme leverages FPGA’s flexibility to optimize model performance by replacing multiplication operations with shift operations, which are less resource-intensive.
+This shift-add structure, as we refer to it, allows for greater parallelism and faster computation, constrained only by the FPGA’s logic resources rather than its limited digital signal processing (DSP) resources[18] on the chip.
+Additionally, we propose a shift parameter quantization technique, where model weights are quantized into integers corresponding to the number of binary shift operations.
+This approach enables real-time processing of the substantial data volumes captured by the DVS system on a mainstream industrial FPGA.
 In a word, the contribution of this work can be listed as follows:
@@
 The results demonstrate comparable performance and generalizability to deep learning models, while achieving real-time recognition over long-distance fibers.
-  • Shallow CNN models, coupled with knowledge distillation, are used to develop a lightweight model that offers high performance and generalizability.
+  • We propose to use shallow CNN models, coupled with knowledge distillation, to develop a lightweight model that offers high performance and generalizability.
 This approach effectively addresses the throughput limitation brought by the high computational complexity of existing algorithms.
-  • A novel hardware acceleration scheme targeted for shallow CNN is developed, leveraging shift-add structures and shift parameter quantization to achieve higher efficiency.
+  • We design a novel hardware acceleration scheme for shallow CNNs, utilizing shift-add structures and shift parameter quantization to significantly improve efficiency.
 II Method
@@
 2.
-The process can be seen as minimizing the difference between the outputs of the student network and the correct results.
-It can be achieved through minimizing a loss function that measures the difference between the outputs of the teacher network and the student network.
 The loss function usually consists of a standard classification loss, like cross-entropy, and a distillation loss that measures the difference between the softened probabilities generated by the teacher and student networks.
@@
 Obviously, both convolution layers and pooling layers only require the input feature map data within the selected window.
-Assume the input to be a spatial-temporal figure, where rows represent discrete times and columns represent positions.
-The data is streamed in row-major order.
+Assume the input is a spatial-temporal figure, where rows represent discrete time points and columns represent spatial positions.
+The data is processed in row-major order.
 Then the activation buffer only needs to store $P$ rows from the input feature map during the computation process.
@@
 The convolution layer only needs to buffer $(P-1-S)*W+Q$ data for each channel instead of the complete input feature map to start operation, where $W$and $S$ denote the width of input feature map and the stride respectively.
-Thank you for the clarification.
-I apologize for the misunderstanding.
-In the case of a convolution layer, it only needs to buffer $(P-1-S)\times W+Q$ data for each channel, rather than the entire input feature map, in order to commence the operation.
-Here, $W$ represents the width of the input feature map, and $S$ represents the stride.
 II-B3 Shift-Add Structure
 The main operation of convolutional neural networks is multiplication and accumulation (MAC).
-Digital signal processing (DSP) block is required to implement MAC units on FPGA.
+DSP block is required to implement MAC units on FPGA.
 The limited number of DSP blocks restricts the number of MAC units that can be implemented on-chip.
@@
 To evaluate and demonstrate the performance of the lightweight model, a 4-layer CNN with a structure outlined in Table III is utilized.
-The model consists of four layers and has a total of 30,771 parameters.
+The model consists of four layers and has a total of 30771 parameters.
 In terms of computational cost, it requires 2,282,496 floating-point operations (FLOPs).
@@
 IV Conclusion
-This paper addresses the throughput limitation issue faced by the DVS systems based on DLMs, which prevents the implementation of real-time recognition over long-distance fibers.
-To overcome this challenge, we propose to use lightweight models as a replacement for DLMs to reduce the computation resource requirements of the system.
-We identify the problem of limited generalizability of the lightweight models and adopts knowledge distillation techniques to enhance the generalizability of the utilized lightweight model.
-Furthermore, we propose an FPGA-based hardware acceleration scheme for fast inference.
-It results in an impressive 2.14x shorter inference time compared to the state-of-the-art GPU implementation, and 4.52x compared to the CPU implementation, which is also the fastest recognition time reported so far in this field.
-Through our proposed method, we achieve real-time recognition over 38-km-long fibers.
-The implemented system allows for real-time recognition over long-distance fibers in practical applications, contributing to the advancement of DVS technology and its use as a smart sensing system in various areas.
+This paper addresses the throughput limitations of DVS systems, which hinder the implementation of real-time recognition over long-distance fibers, which is essential for deploying DVS technology in smart IoT applications.
+To overcome this challenge, we propose using shallow CNN models as a more resource-efficient alternative to DLMs.
+We identify the issue of limited generalizability in lightweight models and address it by applying knowledge distillation techniques to enhance their performance.
+Additionally, we propose a novel hardware acceleration scheme based on FPGAs to accelerate model inference.
+Our approach achieves a significant reduction in inference time, which is 2.14 times faster than state-of-the-art GPU implementations and 4.52 times faster than CPU implementations, making it the fastest recognition method reported in this field to date.
+With this system, we enable real-time recognition over fibers up to 38 kilometers long, advancing the capabilities of DVS technology and its practical application as smart IoT systems.